# Enhancing selectivity using Wasserstein distance based reweighing

### v1

@@ 4 Problem statement and overview of results
2:Output: Compute a distribution $\mathbb{P}_{\mathcal{S}}^{\prime}$ on $\mathcal{S}$ such that the invariant distribution of network weights of a neural net model, trained using SGD with dataset $\mathcal{S}$ and weights $\mathbb{P}_{\mathcal{S}}^{\prime}$, will be closer (in Wasserstein metric) to the invariant distribution of network weights of a neural net model trained using SGD on $\mathcal{T}$ with points weighted as $\mathbb{P}_{\mathcal{T}}$.
3$\triangleright$ RandomSample returns an empirical probability distribution computed from sample size $m$.
4$\triangleright$ $R_{\mathcal{S}}\subseteq\mathcal{S}$ and $R_{\mathcal{T}}\subseteq\mathcal{T}$ denote the random sample of points from their respective ground sets.
5:$\mathbb{P}_{R_{\mathcal{S}}}:=\mathrm{RandomSample}_{m}(\mathcal{S},\mathbb{P}_{\mathcal{S}})$
6:$\mathbb{P}_{R_{\mathcal{T}}}:=\mathrm{RandomSample}_{m}(\mathcal{T},\mathbb{P}_{\mathcal{T}})$
@@
7$\triangleright$ Obtain a $\alpha$-tilted version of $\mathbb{P}_{R_{\mathcal{S}}}$ that’s close to $\mathbb{P}_{R_{\mathcal{T}}}$ using greedy minimum weight metric bipartite matching algorithm (ScaledGreedyReweight in supplement)
8:$\mathbb{P}_{R_{\mathcal{S}}}^{\prime}:=\mathrm{ScaledGreedyReweight}(\mathbb{P}_{R_{\mathcal{S}}},\mathbb{P}_{R_{\mathcal{T}}},\alpha)$
@@
9$\triangleright$ Obtain a reweighted version of $\mathcal{S}$
10:$\mathbb{P}_{\mathcal{S}}^{\prime}=(1-\alpha)\mathbb{P}_{\mathcal{S}}+\alpha\mathbb{P}_{R_{\mathcal{S}}}^{\prime}$.
11$\triangleright$ Train neural net on $\mathbb{P}_{\mathcal{S}}^{\prime}$.
@@
12:Use stochastic gradient descent (SGD) to train the neural net using $\mathbb{P}_{\mathcal{S}}^{\prime}$.
@@
A difference in $\mathbb{P}_{\mathcal{S}}$ and $\mathbb{P}_{\mathcal{T}}$ results in a difference in the convergence point of the weights in any neural net training procedure, like SGD. This is because the loss functions in the SGD algorithm will differ in their weights, though they may have the same form. Therefore, given two mean squared error loss functions weighted with different probability distributions, say $\mathbb{P}_{\mathcal{S}}$ and $\mathbb{P}_{\mathcal{T}}$, on each of their terms, a natural question is: what is the relation between the output of two neural nets that are trained using the two different loss functions?
The expected output of any neural net, on any given input, at the end of a long enough training period, depends on the invariant measure of the weights from the SGD training algorithm. The $1$-Wasserstein distance [^4] See, for example [4] and the references therein for background on the Wasserstein distances. between the weights in the loss functions is the same as the $1$-Wasserstein distance between the occurrence frequencies of data-points in the two underlying data-sets $\mathcal{S}$ and $\mathcal{T}$. Our first theoretical contribution, Theorem 5.4, shows that $W_{1}(\mathbb{P}_{\mathcal{S}},\mathbb{P}_{\mathcal{T}})$, the 1-Wasserstein distance between the loss function weight distributions $\mathbb{P}_{\mathcal{S}}$ and $\mathbb{P}_{\mathcal{T}}$, upper-bounds the total variation distance between the invariant measures of two such neural nets. Therefore, if the data-sets $\mathcal{S}$ and $\mathcal{T}$ are not identical, i.e., $\mathbb{P}_{\mathcal{S}}$ and $\mathbb{P}_{\mathcal{T}}$ do not have the same support set, then the best we can hope for is a small $W_{1}(\mathbb{P}_{\mathcal{S}},\mathbb{P}_{\mathcal{T}})$, in order to obtain similar neural net models for interchangeability. The above explains our choice of the $1$-Wassertein metric in Algorithm 1 (see also Remark 5.5 for comparing with the Levy metric).
@@
Let’s assume $W_{1}(\mathbb{P}_{\mathcal{S}},\mathbb{P}_{\mathcal{T}})$ is large, so that we need to reweigh $\mathcal{S}$. For reweighing, we do not remove any examples from $\mathcal{S}$, instead we prefer to increase the weight of some already present examples in $\mathcal{S}$. Therefore, we compute a distribution $\mathbb{P}_{\mathcal{S}}^{\prime}$ with the same set of support as $\mathbb{P}_{\mathcal{S}}$, such that it minimizes $1$-Wasserstein distance between $(1-\alpha)\mathbb{P}_{\mathcal{S}}+\alpha\mathbb{P}_{\mathcal{T}}$ and $\mathbb{P}_{\mathcal{S}}^{\prime}$, for some fixed choice of tilt factor $\alpha\in[0,1]$.555 The optimum value of $\alpha$ can be chosen by binary search after running multiple training evaluations. We use $\mathbb{P}_{\mathcal{S}}^{\prime}$ as the new set of weights for neural net training. Note that, for $\alpha$ close to $1$, the distribution will be closest to $\mathbb{P}_{\mathcal{T}}$ while being supported on $\mathcal{S}$.
@@
Note that the optimal $\mathbb{P}_{S}^{\prime}$ mentioned above can be computed by solving a linear program that closely resembles the $1$-Wasserstein distance computation linear program. However, the size of the linear program would be quadratic in the size of the data-sets, making the computation intractable for most practical data-sets. [^6] A typical large data-set has 10-100M examples, and computing $W_{1}$ over two such data-sets requires solving a linear program – a $\Theta(n^{3})$ time procedure, resulting in the order of $10^{24}$ computational operations! Therefore, we look for inaccurate but efficient algorithms and a natural candidate is the randomized greedy algorithm below.
The 1-Wasserstein metric has an equivalent interpretation as a optimal transport problem. See Theorem 5.6 (essentially repeated from [23]) for a formal statement that reduces it to the metric minimum weight bipartite matching problem [20].
@@
One tractable way to compute a minimum weight bipartite matching is to use a faster but sub-optimal algorithm. The greedy algorithm, formally studied by [20] in this context, is a natural contender. It is almost linear time, and easy to implement. However, [20] showed that such greedy algorithms can be really inaccurate. Moreover, even the greedy algorithm requires linear space and given the size of our data-sets, that can also become a constraint.
However, if our input instances admit a small sized covering then we show that the greedy algorithm run on a large enough random sample of data, i.e., Algorithm 1 for a large enough choice of $m$, performs reasonably well. Our main contribution here is Theorem 5.16. Theorem 5.16 states that the greedy algorithm on a small fraction random sample of an $\Theta(n)$ point data-set can be used to approximate the $1$-Wasserstein distance with a poly-logarithmic factor approximation.
So, the question arises: What precisely does a small covering assumption above mean and what for kind of natural problems does Algorithm 1 scales efficiently without deterioration in the approximation guarantee?
The metric entropy of a point set (see Definition 5.10) is the minimum number of balls of a given radius required to cover the point set. So, in high dimensions, data-sets with low metric entropy can be characterized from their values on (relatively) small balls spread through space. This is indeed the case with DEL data-sets like MNK1-MNK2 in our drug discovery application. The combinatorial synthesis process utilized in DELs often results in local chemical similarity among compounds that share common building blocks. Since similar molecules likely have the same binding behavior, synthesized molecules form a small ball around a parent molecule in the molecule fingerprint space. Therefore, molecule binding vs non-binding data-sets likely have low metric entropy. It turns out that for training data-sets with low metric entropy the greedy algorithm of [20] performs provably well (cf. Theorem 5.13).
The proof of Theorem 5.16 relies on Theorems 5.13 and 5.15.
Theorem 5.13 show that greedy minimum weight matching on bipartite graphs for vertex sets with low metric entropy has a much better (poly-logarithmic) approximation guarantee in our case, as opposed to the polynomial approximation guarantee from [20]. This argument, especially the connection between covering and matching in Lemma 5.12 may be of independent interest.
@@
Theorem 5.15 essentially shows that random samples on data-sets with bounded metric entropy preserves $1$-Wasserstein distances. This allows us to work with small samples of large data-sets. For the proof, we need large deviation bounds for the $1$-Wasserstein distance between the theoretical distribution and its empirical distribution. Such results have been explored previously with tight Sanov’s theorem type bounds in low dimensional spaces (see for example [4]), but in high dimensions, we need the assumption of low metric entropy for the same results to go through (see chapter 6 in [9]). Coincidentally, that is precisely our assumption in Theorem 5.13!
5 Theorems and Proofs
5.1 Bounding $1$-Wasserstein distance suffices

### v2

@@ 3.2 Reweighing algorithm
2:  Output: Compute a distribution $\mathbb{P}_{\mathcal{S}}^{\prime}$ on $\mathcal{S}$ such that the invariant distribution of network weights of a neural net model, trained using SGD with dataset $\mathcal{S}$ and weights $\mathbb{P}_{\mathcal{S}}^{\prime}$, will be closer (in Wasserstein metric) to the invariant distribution of network weights of a neural net model trained using SGD on $\mathcal{T}$ with points weighted as $\mathbb{P}_{\mathcal{T}}$$\triangleright$ Algorithm starts:$\triangleright$ RandomSample returns an empirical probability distribution computed from sample size $m$$\triangleright$ $R_{\mathcal{S}}\subseteq\mathcal{S}$ and $R_{\mathcal{T}}\subseteq\mathcal{T}$ denote the random sample of points from their respective ground sets.
3:  $\mathbb{P}_{R_{\mathcal{S}}}:=\mathrm{RandomSample}_{m}(\mathcal{S},\mathbb{P}_{\mathcal{S}})$
@@
4:  $\mathbb{P}_{R_{\mathcal{T}}}:=\mathrm{RandomSample}_{m}(\mathcal{T},\mathbb{P}_{\mathcal{T}})$$\triangleright$ Obtain a $\alpha$-tilted version of $\mathbb{P}_{R_{\mathcal{S}}}$ that’s close to $\mathbb{P}_{R_{\mathcal{T}}}$ using greedy minimum weight metric bipartite matching algorithm, described as GreedyAlgorithm (Algorithm 2) in supplement
@@
5:  $\mathbb{P}_{R_{\mathcal{S}}}^{\prime}:=\mathrm{GreedyAlgorithm}(\mathbb{P}_{R_{\mathcal{S}}},\mathbb{P}_{R_{\mathcal{T}}},\alpha)$$\triangleright$ Obtain a reweighted version of $\mathcal{S}$
6:  $\mathbb{P}_{\mathcal{S}}^{\prime}=(1-\alpha)\mathbb{P}_{\mathcal{S}}+\alpha\mathbb{P}_{R_{\mathcal{S}}}^{\prime}$$\triangleright$ Train neural net on $\mathbb{P}_{\mathcal{S}}^{\prime}$.
@@
7:  Use stochastic gradient descent (SGD) to train the neural net using $\mathbb{P}_{\mathcal{S}}^{\prime}$.
3.3 Theoretical results
The rest of this section concentrates on providing a theoretical explanation for why Algorithm 1 should work as intended and scale computationally.
3.3.1 Choice of $1$-Wasserstein metric
@@
In our problem statement, a difference in $\mathbb{P}_{\mathcal{S}}$ and $\mathbb{P}_{\mathcal{T}}$ results in a difference in the convergence point of the weights in any neural net training procedure, like SGD. This is because the loss functions in the SGD algorithm will differ in the weights of their summand terms, even though they may have the same form. Therefore, given two mean squared error loss functions weighted with different probability distributions, say $\mathbb{P}_{\mathcal{S}}$ and $\mathbb{P}_{\mathcal{T}}$, on each of their terms, a natural question is: what is the relation between the limiting distribution of network weight parameters of two neural nets that are trained using the two differently weighed loss functions?
Our first theoretical contribution, Theorem 3.2, shows that $W_{1}(\mathbb{P}_{\mathcal{S}},\mathbb{P}_{\mathcal{T}})$, the 1-Wasserstein distance between the loss function weight distributions $\mathbb{P}_{\mathcal{S}}$ and $\mathbb{P}_{\mathcal{T}}$, upper-bounds the total variation distance between the invariant measures of two such neural nets under a covariate shift like assumption. Note that the covariate shift assumption is used in distribution shift correction literature, see for example Bickel et al. (2009). More formally:
Assumption 3.1.
Let $f(w,x)$ be the neural network output, for weights $w$, input $x$ and corresponding label $y$. We assume that $x\in Q_{d}$ (the $d$ dimensional boolean hypercube), $f$ and $y$ are bounded, $y,f(w,x)\in[0,1]$ and for all $w$:
Assumption 3.1 is weaker than the usual covariate shift assumption, since in the latter case, the RHS of Equation 4 would equal 00.888The assumption can be further weakened by not requiring that it needs to hold for all $w$, but only $w$ near local optima; see the discussion in the supplement Section B.1. Under Assumption 3.1, we show the following theorem:
Theorem 3.2.
(see Theorem B.3 for precise statement) Suppose we train two neural networks, such that (1) the limiting stochastic differential equation (SDE) corresponding to the training SGD (as SGD step-size goes to 00) is strongly elliptic, [^9] This ensures the invariant measure of the SDE exists, is smooth and unique. and (2) Assumption 3.1 holds, on different input distributions, $\mathbb{P}_{\mathcal{T}}$ and $\mathbb{P}_{\mathcal{S}}$, using the stochastic gradient descent (SGD) algorithm. Then, if $W_{1}(\mathbb{P}_{\mathcal{S}},\mathbb{P}_{\mathcal{T}})=\Omega(1)$ (the interesting case of our problem), the total variation distance between their invariant measures can be bounded by $O(W_{1}(\mathbb{P}_{\mathcal{T}},\mathbb{P}_{\mathcal{S}}))$, for the limiting SDE of the SGD.
Therefore, the above explains our choice of the $1$-Wassertein metric as the metric to use in the greedy minimum weight metric bipartite matching computation in Algorithm 1.
3.3.2 Metric bipartite matching
@@
The next question is, suppose we want to compute a distribution $\mathbb{P}_{\mathcal{S}}^{\prime}$ with set of support $\mathcal{S}$, such that it minimizes $1$-Wasserstein distance between $(1-\alpha)\mathbb{P}_{\mathcal{S}}+\alpha\mathbb{P}_{\mathcal{T}}$ and $\mathbb{P}_{\mathcal{S}}^{\prime}$, for some fixed choice of tilt factor $\alpha\in[0,1]$.101010 The optimum value of $\alpha$ can be chosen by trial and error after running multiple training and validations, to reduce any over-fitting. We would then use $\mathbb{P}_{\mathcal{S}}^{\prime}$ as the new set of weights for neural net training.
@@
While the optimal $\mathbb{P}_{\mathcal{S}}^{\prime}$ mentioned above can be computed by solving a linear program (LP) that closely resembles the $1$-Wasserstein distance computation LP, the number of constraints would be quadratic in the size of the data-sets, making the computation intractable for large datasets. [^11] A typical large data-set has 10-100M examples, and computing $W_{1}$ over two such data-sets requires solving a linear program – a $\Theta(n^{3})$ time procedure, resulting in the order of $10^{24}$ computational operations! Therefore, we look for inaccurate but efficient algorithms and a natural candidate is the randomized greedy algorithm for minimum weight metric bipartite matching, for reasons explained below.
@@
The 1-Wasserstein metric has an equivalent interpretation as an optimal transport problem. In fact, it is equivalent to solving the metric minimum weight bipartite matching problem Reingold and Tarjan (1981). The reduction is fairly intuitive and consists of duplicating the supply and demand points, in the optimal transport problem, in proportion to their weights in the dataset. Theorem B.5 (essentially repeated from Sharathkumar and Agarwal (2012)) provides a formal statement reducing the former to the latter.
One tractable way to compute a minimum weight bipartite matching is to use a faster but sub-optimal algorithm. The greedy algorithm, formally studied by Reingold and Tarjan (1981) in this context, is a natural contender. In this paper, we show that if our input instances admit a small sized covering then the greedy algorithm, i.e., Algorithm 1, provides a better approximation guarantee than the worst case lower bound from Reingold and Tarjan (1981). Our main contribution here is Theorem 3.3, which is just a conjunction of Theorems 3.4 and 3.5. Theorem 3.4 provides the combinatorial argument around the for the greedy algorithm under the small covering assumption, and is discussed in Subsection 3.3.3. Theorem 3.3 can be informally stated as follows.
Theorem 3.3.
(see Theorem B.16 for general statement with trade-offs) Suppose we are given two data-sets with $\mathcal{S}$ and $\mathcal{T}$ that are weighted according to distributions $\mathbb{P}_{\mathcal{S}}$ and $\mathbb{P}_{\mathcal{T}}$. If,
  1. 1.
    Small covering: $\mathcal{S}\cup\mathcal{T}$ admits a covering with $\eta$ $\ell_{1}$-balls of radius $\zeta$; with $\eta,\zeta=O(\log^{c}n)$, and $c\leq\frac{1}{2(1+\log_{2}(3/2))}$; and
  2. 2.
    $\mathbb{P}_{\mathcal{S}}$ and $\mathbb{P}_{\mathcal{T}}$ are sufficiently far apart: $W_{1}(\mathbb{P}_{\mathcal{S}},\mathbb{P}_{\mathcal{T}})\geq\log\log n+S_{n}$, where $S_{n}$ measures how spread out the the covering balls are (see Definition B.14).
then the greedy algorithm achieves an approximation ratio of $O(d^{c^{\prime}})$ for some $c^{\prime}\leq 0.73$, [^12] The value $0.73$ comes from using $\xi=2$ in the bound in Theorem 3.4. with probability $1-o(1)$, when computed on a small random sample of $r(n)$ fraction of data-points and $r(n)\to 0$.
3.3.3 Small covering assumption
So, the question arises: What does a small covering assumption above mean in the context of minimum weight metric bipartite matching algorithms, and what is its underlying combinatorial connection to such matchings?
One way to specify small coverings is via metric entropy – the minimum number of balls of a given radius required to cover the point set (see Definition B.9). It turns out that for computing minimum weight bipartite matchings on pointsets with low metric entropy, the greedy algorithm of Reingold and Tarjan (1981) performs provably well.
Theorem 3.4.
(see Theorem B.13) For $d=\Theta(\log n)$, let $\eta=O(d^{\frac{1}{\xi\log_{2}3/2}})$ and $\zeta=O(\eta)$. Suppose that the input points can be covered by $\eta$ $\ell_{1}$-balls of radius $\zeta$, then the greedy algorithm achieves an approximation factor of $\max\{2\zeta,O\left(d^{\frac{1+\xi\log_{2}(3/2)}{\xi(1+\log_{2}(3/2))}}\right)\}$ for $\xi>1$, on points on the $d$ dimensional hypercube.
Note that for points on the $d$ dimensional hypercube, an approximation factor of $d$ for minium weight metric bipartite matching is trivial, but we improve it to $O(d^{c})$ for $c<1$, exact $c$ depends on the metric entropy of the dataset (see Corollary B.12 and Theorem B.13 for precise trade-offs).
The crux of the proof of Theorem 3.4 consists of a structural characterization of alternating cycles [^13] An alternating cycle in a matching is simply a cycle consisting of alternate matched and unmatched edges. and matchings (Lemma B.11) that may be of independent interest. The key to the proof of Lemma B.11 is the following (informal) idea: Let $\gamma$ be an alternating cycle in the greedy matching. Suppose that the sum of weights of the matched edges between vertices in $\gamma$ is $\kappa$ times their weight in the minimum weight matching, i.e., the cycle $\gamma$ is long. But, how can a cycle be long in a metric space like the hypercube, which has diameter $\log n$, and number of vertices $n$? The answer is that the cycle $\gamma$ must have many long edges. That together with the assumption that there exists a small covering implies a contradiction for an appropriate choice of $\kappa$.
3.3.4 Greedy on random sample
Algorithm 1 uses the greedy algorithm on top of a small random sample of the datasets to deal with the quadratic time complexity when datasets $\mathcal{S}$ and $\mathcal{T}$ are large. Therefore, we show that using a small random sample does not lead to large deterioration in the approximation guarantee. Theorem 3.5 states that using random samples for datasets with bounded metric entropy do not lead to a much worse approximation guarantee, if the $W_{1}$ distance between the empirical distribution computed from $m$ samples ($\hat{\mu}_{m}$) and the true distribution ($\mu$) is known to be not too small (which is the interesting case of the problem).
Theorem 3.5.
(Informal; see Theorem B.15 for precise formal statement) For a dataset with $\log^{O(1)}n$ metric entropy with $\ell_{1}$ balls of $\log^{O(1)}n$ radius, and a random sample of size $m=o(n)$, the 1-Wasserstein distance between the empirical distribution and the true distribution of data-sets with bounded metric entropy obeys the following Sanov type concentration bound:
${\scriptstyle\exists\ m=o(n),\ \lim_{n\to\infty}\frac{1}{m}\ln\mathbb{P}\left(W_{1}(\hat{\mu}_{m},\mu)\geq\log\log n+S_{n}\right)\leq-\Omega(1),}$ (2)
where $S_{n}$ and it measures how spread out the the covering balls are (see Definition B.14).
Note that, $S_{n}\leq\log n$, for the hypercube, so when $S_{n}=o(\log n)$ but the diameter of our point set is $O(\log n)$, i.e., most points are clustered around some $x_{0}$ except for small fraction of outliers, then one can use a random sample to reduce the input size further without deterioration of the worst case approximation factor.
@@ 3.3.4 Greedy on random sample
For the proof of Theorem 3.5, we need large deviation bounds for the $1$-Wasserstein distance between the theoretical distribution and its empirical distribution. Such results have been explored previously with tight Sanov’s theorem type bounds in low dimensional spaces (see for example Bolley et al. (2007)). However, our underlying space has large dimension, i.e., $\log n$, which depends upon $n$. The constants in the exponential in the theorems in Bolley et al. (2007) will thus depend on $n$, and it’s not immediately clear to us whether the dependence can be easily removed. Hence, we need the assumption of low metric entropy for the same results to go through (see chapter 6 in Dembo and Zeitouni (2010)).
4 Example application: drug discovery
A natural question is, when does the small covering assumption hold in practice? This seems to happen in the drug discovery setting. [^14] The combinatorial synthesis process utilized in DNA encoded library (DEL) compounds often results in local chemical similarity among compounds that share common building blocks. Since similar molecules likely have the same binding behavior, synthesized molecules form a small ball around a parent molecule in the molecule fingerprint space. Therefore, molecule binding vs non-binding data-sets likely have low metric entropy. So, as a concrete motivating example, we illustrate an application of Algorithm 1 to a toy problem in the drug discovery area (see also Section D).
In drug discovery, typically, one wants to isolate selective small molecules (inhibitors) that bind strongly to a given enzyme, but often we want to exclude small molecules that bind to another similar enzyme. For example, MNK1 and MNK2 are two structurally similar kinases (a kinase is an enzyme for phosphorylation or de-phosphorylation of proteins). We want to identify small molecules that bind strongly to MNK2 (MNK2 hits), but we also prefer that the identified small molecules not bind to MNK1 (MNK1 non-hits). In other words, we want to isolate molecules that are selective for MNK2 over MNK1.
Performance improvements: In the in-silico experiments, we were able to increase the percentage of MNK1 non-hits in our set of top predicted MNK2 hits – the selectivity – from 54% to 95% on holdout data, using the reweighing procedure in Algorithm 1. We are not aware of other such multi-target prediction results in DNA encoded library (DEL) space (see Satz et al. (2022) for background), where one simultaneously predicts hits/non-hits against two or more proteins. However, the success rates for single target experiments with traditional high-throughput screening is $\sim 1\%$ (see for example the discussion in McCloskey et al. (2020)) and it is generally accepted that multi-target prediction is a harder problem.
Training: We used a relatively small training set of about $250$K small molecules in total; labeled as MNK1 non-binders, and MNK2 binders as well as non-binders. To evaluate the effect of reweighting on the selectivity [^15] Recall that, we want to find small molecules that are MNK2 binders but preferrably MNK1 non-binders. of predicted MNK2 binders, we use a small holdout set of $7$K small molecules that consists of molecules which are labeled as: MNK2 hits (binders), and MNK1 hits (binders) or MNK1 non-hits (non-binders). In Figure 1, for the neural network models with and without reweighing, we plot the cumulative number of MNK1 non-hits on the $y$-axis; and on the $x$-axis any given point, say $k$, represents the top $k$ predicted MNK2 hits from the examples in the holdout set.
While we can not make our training data-sets and code public for proprietary reasons, we were able to experimentally (in wet-lab) verify that two out of the top fifty (actually 43, since 7 out of 50 molecules could not be synthesized and tested) predicted selective small molecules, obtained by running our neural network model on the Enamine 1.9B molecules catalog (https://enamine.net), were indeed selective for MNK2 over MNK1.161616The compounds are Z1918489591 and Z5890616727 in the enamine catalog. That is a success rate of roughly 5% on this admittedly small sample set. We do note that the results are from a single point concentration assay and can be noisy.
Assay based validation: More importantly, the two predicted and assay tested molecules in Figure 2 provide a degree of verification for our experimental application (which has been the motivation, but is not the focus of this paper).
Effect of reweighing: The main aim of the following discussion is to justify Remark 4.1 below.
Remark 4.1.
Algorithm 1 is successfully highlighting a closely packed portion of the chemical space of MNK2 binders that are non-binders to MNK1, while the top predicted MNK2 binders in the baseline model are relatively more spread out in the chemical fingerprint space.
In order to observe the effects of reweighing on the top $100$ predicted binders to MNK2, we did four types of similarity comparisons in between two sets: (1) top $100$ baseline predictions, and (2) top $100$ treatment (reweighed) predictions: [^17] Recall that, the baseline training dataset is not weighed using MNK1 data, so it effectively just predicts small molecules that are potent MNK2 binders without regard to their MNK1 binding property. The experiment training dataset reweighs the MNK2 binders and non-binders to highlight MNK1 non-binders. Therefore, if the reweighting is done optimally, with just the right amount of tilt, then we may be able to predict potent MNK2 binders that are preferrably MNK1 non-binders.
  1. 1.
    For every baseline molecule we computed its mean similarity with the remaining $99$ baseline molecules,
  2. 2.
    For every treatment (reweighed) molecule we computed its mean similarity with the remaining $99$ treatment molecules.
  3. 3.
    For every baseline molecule we computed its mean similarity with the $100$ top treatment molecules,
  4. 4.
    For every treatment molecule we computed its mean similarity with the $100$ top baseline molecules.
We expect the top predictions to qualitatively have the following properties:
  • The treatment (reweighing) highlights a small portion of the molecular space of MNK2 binders the ones that are not MNK1 binders, so we expect the mean similarity in (1) to be smaller than (2).
  • If we assume there’s a small region of molecules in the fingerprint space which are selective, i.e., bind to MNK2 but not MNK1, and from which the baseline has sampled about $50\%$ of its top $100$ predictions, while the treatment has sampled about $95\%$ of its top $100$ predictions (see Figure 1); then we expect the following: (1) a plot of the computed similarities in item (3) above will have a bimodal distribution with about $50\%$ of the mass in each mode, and (2) a plot of the computed similarities in item (4) will have a bimodal distribution with about $95\%$ of the mass in one mode.
Indeed our experiments are consistent with the above. In our plots, we found that the top 100 predicted binders in the experimental model are indeed more similar to each other than the baseline model (mean Tanimoto similarity increases from $0.32$ to $0.53$), and hence are ”packed” more closely together (see Figure 3).
Figure 3: (L) Tanimoto similarities of top molecules in base model vs (R) Tanimoto similarities of top molecules in reweighed model
We also observe that the mean similarity score for the cross comparison, in Figure 4(L), is clearly bifurcated into two parts with means around $0.2$ and $0.5$ respectively. The right peak of 30 small molecules mostly comes from the 54 molecules that were MNK1 non-binders. This means that about half of the top 100 predicted MNK2 binders in baseline are much more similar to the top 100 predicted MNK2 binders in the reweighed model. Thus, Figures 3 and 4 are consistent with Remark 4.1.
Figure 4: (L) Mean Tanimoto Similarity for each top molecule in base with reweighed model (note the bifurcation) vs (R) Mean Tanimoto Similarity for each top molecule in reweighed with base model.
5 Acknowledgements
Appendix A Supplement
In this appendix, we first formally state the assumptions and outline our theorems in Section B, and then provide the full proofs in Section C. Section D contains some more details about the drug discovery application.
Appendix B Formal setup and theorem statements
B.1 Choice of metric in Algorithm 1: Bounding $1$-Wasserstein distance suffices